OpenStudy (andu1854):

Find all points (x,y) on the graph of y = x/(x-2) with tangent lines perpendicular to the line y=2x-1.

OpenStudy (mrnood):

ok - so first: what is the slope of y=2x-1? then what is the slope of a line perpendicular to it?

OpenStudy (andu1854):

the slpe of y = 2x-1 would be m =2 and for a perpendicular line, the slope would be m =-1/2

OpenStudy (mrnood):

ok Now the tangent line of any curve has the slope of the derivative at that point so you need to get an equation for the derivative of y = x/(x-2) (use the 'quotient rule')

OpenStudy (andu1854):

so the derivative I got was -2/(x-2)^2

OpenStudy (mrnood):

I haven't checked that - but assume it OK So the points where the derivative are perp to the line are where the derivitiv = -1/2 SO make that equation and solve it

OpenStudy (mrnood):

I'll check the derivitive

OpenStudy (andu1854):

-2/(x-2)^2 = -1/2?

OpenStudy (mrnood):

yes, if the derivative is correct

OpenStudy (andu1854):

ok so I got x = 4,0

OpenStudy (mrnood):

f'(u/v) = (u dv/dx - v du/dx)/ v^2

OpenStudy (andu1854):

I meant -4

OpenStudy (mrnood):

OK - I checked your differentiation and it is correct so you have your answers - well done

OpenStudy (mrnood):

I think it IS 0, 4 not 0,-4

OpenStudy (andu1854):

yeah, you are correct

OpenStudy (mrnood):

So - you did ALL that without any problem - what was your initial difficulty?

OpenStudy (andu1854):

I was using the wrong slop, parallel slope instead of the perpendicular slope and I just wanted to also make sure I was doing every thing else correct...

OpenStudy (mrnood):

oops wait a moment I need to check - I think I missed a - sign...

OpenStudy (andu1854):

I was setting the derivative equal to -2 instead of -1/2, but I got the right answer so you were a very big help... thank you once again...

OpenStudy (mrnood):

OK - if you're happy, I'll leave it as it is

OpenStudy (mrnood):

OK - I see where I thought I'd gone wrong - but it was OK The above answers are correct